# Phase driven exact and unconventional superradiance phase transition in non-Hermitian cascaded Rabi cavities

### v1

@@ IV Signatures of the superradiance phase transition
We detect the signatures and check the phase boundary of the SPT. We start by studying the ground-state energy to reveal the SPT. The energies are sorted by the real parts of the energies, and the ground-state energy means the energy with the lowest real part. Under $\varphi=\frac{\pi}{2}$ and different $\eta$, the real parts of the ground-state energies ${\rm Re}(E_{0})/\eta$ as a function of $\mathcal{J}$ for $g=0.7$ are plotted in Fig. 2(a). The photonic truncation number is taken at $N=80$. As can be seen that with the emergence of SPT, ${\rm Re}(E_{0})/\eta$ for different $\eta$ decrease from $-1$ to lower values. According to the obtained analytical result, the transition point under $\varphi=\frac{\pi}{2}$ and $g=0.7$ should be at $\mathcal{J}=0.745$. Intuitively, the SPT occurs at the transition points predicted by the exactly analytical result (shown with the vertical black line). Meanwhile, we calculate ${\rm Re}(E_{0})/\eta$ in the $\varphi=\frac{\pi}{4}$ case and ${\rm Re}(E_{0})/\eta$-$\mathcal{J}$ relations for $g=0.4$ are plotted in Fig. 2(b). Similar to the $\varphi=\frac{\pi}{2}$, when $\mathcal{J}$ exceeds to the transition point, ${\rm Re}(E_{0})/\eta$ for different $\eta$ in the $\varphi=\frac{\pi}{4}$ case decrease from $-1$ to lower values, signaling the appearance of SPT as well. From the analytical expression of phase boundary, it is known that the transition point under $\varphi=\frac{\pi}{2}$ should be at $J=0.5$ regardless of what value of $g$ is. It is readily seen that for $g=0.4$, the transition points of ${\rm Re}(E_{0})/\eta$ occur at the analytically obtained transition points. Although the ground state energy of the system decreases after the occurrence of conventional and unconventional SPTs, it should be pointed out that they do not belong to the same phase transition type. As shown in Fig. 1(c), the second-order derivative of ${\rm Re}(E_{0})/\eta$ under $\varphi=\frac{\pi}{2}$ and $g=0.7$ is not continuous at the the transition point. Figure 1(d) shows that the first-order derivative of ${\rm Re}(E_{0})/\eta$ under $\varphi=\frac{\pi}{4}$ and $g=0.4$. The results show that the conventional SPT should be the second-order quantum phase transition, while the unconventional SPT should be the first-order quantum phase transition. In fact, only the $|\varphi|=\frac{\pi}{4}$ and $|\varphi|=\frac{3\pi}{4}$ cases correspond to the second-order quantum phase transition, while the cases with other $\varphi$ correspond to the first-order quantum phase transition. However, we are more concerned with conventional and unconventional superradiance phase transitions in this paper, so the results of other $\varphi$ are not considered
@@
A reasonable explaination for the macroscopic photon excitations in the suprradiance phase is that the atoms are excited and become active. To check this point, the atom populations $\langle\sigma^{1(2)}_{z}\rangle$ as a function of $\mathcal{J}$ for different $\varphi$ and $g$ are plotted in Fig. 4(a)-4(d). We can see that when $\mathcal{J}$ is smaller than the transition point, the atoms are stable in the ground state, with populations $\langle\sigma^{1(2)}_{z}\rangle=-1$. When $\mathcal{J}$ crosses the transition point, it is seen that the populations of atoms deviate from -1, implying that the atoms are excited and become active. The excited atoms dominate the macroscopic photon excitations in the SPT. Although the photon excitations and the atom populations of the system vary obviously with occurrence of SPT, the two types of SPTs lead to different varying behaviors of observables. From Fig. 3, we can see that when the conventional SPT happens, $\langle a^{{\dagger}}_{1(2)}a_{1(2)}\rangle$ continuously varies to macroscopic value, while the unconventional SPT happens, $\langle a^{{\dagger}}_{1(2)}a_{1(2)}\rangle$ directly jumps to macroscopic values without asymptotic process. Similarly, we can see the obvious difference between the two types of SPTs from the atomic populations in Fig. 4.
@@
In addition to the photon population and the atom population, the excitation gaps behind the two types of SPTs are different as well. The excitation gaps here contain two meanings. One is the real excitation gap ${\rm Re}(\Delta_{E})/\eta$ between the real parts of the lowest two energies. Another is the imaginary excitation gap ${\rm Im}(\Delta_{E})/\eta$ between the imaginary parts of the lowest two energies. Figures. 5(a1)-5(d1) (Figs. 5(a2)-5(d2)) presents ${\rm Re}(\Delta_{E})/\eta$ (${\rm Im}(\Delta_{E})/\eta$) as a function of $\mathcal{J}$ under different $\eta$. We can see that no matter $\varphi=\frac{\pi}{2}$ and $\varphi=\frac{\pi}{4}$, the real excitation gap experiences a gap closing process when $\mathcal{J}$ exceeds the transition point. However, for the imaginary excitation gap, there is an obvious difference between the two studied cases. From Figs. 5(a2) and 5(b2), we can see that ${\rm Im}(\Delta_{E})/\eta$ always keeps zero (${\rm Im}(\Delta_{E})/\eta$ far less than $10^{-14}$ can be numerically regarded as zero.) before and after the transition point. From Figs. 5(c2) and 5(d2), we can see that the imaginary excitation gaps experience the gap closing process, similar to the circumstances for the real excitation gaps.
@@
It is noted that for a $\mathcal{PT}$-symmetric Rabi system, there exists real-complex transition in energy spectrum PT ; PT1 ; es_NH_2 ; es_NH_3 ; es_NH_4 . Surprisingly, for the non-Hermitian two-Rabi cavity system without $\mathcal{PT}$ symmetry, there exists real energy spectrum as well. Under $\varphi=\frac{\pi}{2}$, the maximal imaginary of the energy spectra for $g=0.7$ and $g=0.8$, i.e., ${\rm Im}(E)_{max}$ are plotted in Figs. 6(a) and 6(b), respectively. The analytically obtained superradiance transition point for $g=0.7$ is $J=0.745$, and that for $g=0.8$ is $J=0.82$. ${\rm Im}(E)_{max}/\eta$ approaches to zero before and after the superradiance transition point implying that the energy spectra under $\varphi=\frac{\pi}{2}$ are real, and there is no real-complex transition in energy spectrum. Compared to the $\varphi=\frac{\pi}{2}$ case, the spectral property of the $\varphi=\frac{\pi}{4}$ is different. Under $\varphi=\frac{\pi}{4}$, ${\rm Im}(E)_{max}/\eta$ for $g=0.4$ and $g=0.6$ are plotted in Figs. 6(c) and 6(d), respectively. Different from the $\varphi=\frac{\pi}{2}$ cases, here the energy spectra are complex because of the nonvanishing ${\rm Im}(E)_{max}/\eta$. In fact, the real and complex characteristics of the energy spectrum can be understood analytically as well. From Eq. (10) and Eq. (13), we know that when $\varphi=\frac{\pi}{2}$, the imaginary part disappears, but when $\varphi=\frac{\pi}{4}$, the imaginary part still exists. The results in Fig. 6 provide numerical validation of the previous analytical derivation. Considering the fact that the $\varphi=\frac{\pi}{2}$ cases have full real energy spectra, we conjecture that here the conventional SPT belongs to the same universality class as the Hermitian Rabi systems HP ; ps2 ; qc_1 ; qc_2 ; qc_3 ; zhangyy_1 and the critical exponent of the phase transition can be obtained from the critical behavior. However, for the complex energy spectrum case ($\varphi=\frac{\pi}{4}$), we conjecture that there is no critical behavior. Although the existence of quantum criticality of conventional SPT and the absence of quantum criticality of unconventional SPT have been reflected in the photon population and atom population, we still want to further verify them by studying the ground-state fidelity and fidelity susceptibility.
@@
V conventional and unconventional superradiance phase transition
We focus on the quantum phase transition and its critical behavior triggered by the cavity coupling parameter $\mathcal{J}$. The fidelity for the ground state $\mathcal{F}$ is defined as fidelity_0 ; fidelity_1 ; fidelity_2 ; fidelity_3
$\mathcal{F}=|\langle\psi_{0}(\mathcal{J})|\psi_{0}(\mathcal{J}+\delta\mathcal{J})|,$ (14)
which weighs the survival probability of the ground state after the cavity coupling parameter $\mathcal{J}$ from an infinitesimal distance $\delta\mathcal{J}$. Further, according to Refs. FS_1 ; FS_2 ; FS_3 ; FS_4 ; FS_5 , we can obtain the fidelity susceptibility $\chi_{\mathcal{F}}$ as
@@
If there exists quantum phase transition, the fidelity will experience a drop at the phase transition point, and the fidelity susceptibility will peak at the phase transition point.
@@
Under $\varphi=\frac{\pi}{2}$, $\chi_{\mathcal{F}}/\eta$ for $g=0.7$ and $g=0.8$ are plotted in Figs. 7(a) and 7(b), respectively. As seen that with the increase of $\mathcal{J}$, $\chi_{\mathcal{F}}/\eta$ smoothly reaches the peak and then smoothly decreases. Besides, $\chi_{\mathcal{F}}/\eta$ presents obvious critical behaviors, because it is seen that with the increase of $\eta$, the peak of $\chi_{\mathcal{F}}$ at the coupling parameter $\mathcal{J}_{max}$ gradually tends to the exact solution. Moreover, with the increase of $\eta$, the peak of $\chi_{\mathcal{F}}$ get higher. However, for the $\varphi=\frac{\pi}{4}$ cases, there is no critical phenomenon. Under $\varphi=\frac{\pi}{4}$, the fidelity $\mathcal{F}$ for $g=0.4$ and $g=0.6$ are plotted in Figs. 7(c) and 7(d), respectively. Intuitively, $\mathcal{F}$ sharply drop to zero at single parameter $\mathcal{J}$ and there is no asymptotic behavior between the valley of $\mathcal{F}$ (or the peak of $\chi_{\mathcal{F}}$) and $\eta$. In other words, the conventional SPT exists quantum criticality while there is no quantum criticality for the unconventional SPT.
Next, we study the quantum criticality of the conventional SPT ($\varphi=\frac{\pi}{2}$ case). From the $\chi_{\mathcal{F}}$-$\eta$ curves, the maximal fidelity susceptibility $\chi^{max}_{F}$ under a specific $\eta$ is obtained. Then, the universal finite $\eta$ scaling relation of $\chi_{\mathcal{F}}$ for $g=0.7$ is plotted in Fig. 8(a). The black fitting curve satisfies $\ln\chi_{\mathcal{F}}=(-1.363\pm 0.004)\ln\eta+2.388$. It implies the scaling exponent $\mu$ of $\chi_{\mathcal{F}}$ is $\mu=-1.363\pm 0.004$. As introduced in Ref. qc_1 ; RMP2011 , the relationship between $\mu$ and $\nu$ obeys $\mu=\frac{2}{\nu}$, by which we obtain $\nu=1.4674\pm 0.004$. Although the numerically obtained critical exponents $\mu$ and $\nu$ has slightly numerical deviation from the exact solution (with $\nu_{exact}=3/2$) of the single Rabi model HP ; ps2 , we argue that the quantum phase transition in the $\varphi=\frac{\pi}{2}$ is the same universality class as that in the single Rabi model. Within the neighborhood of the phase transition point, $\chi_{\mathcal{F}}$ scales as $\left[\chi^{max}_{\mathcal{F}}-\chi_{\mathcal{F}}\right]/\chi_{\mathcal{F}}=f\left[\eta^{1/\nu}\left(\mathcal{J}-\mathcal{J}_{max}\right)\right]$, where $f(x)$ is a universal scaling function. Under different $\eta$, we find the corresponding data of $\left[\chi^{max}_{\mathcal{F}}-\chi_{\mathcal{F}}\right]/\chi_{\mathcal{F}}$ are collapsed into one curve with the exact critical exponent $\nu=\frac{3}{2}$, signaling the same universal class as the single Rabi model HP ; ps2 . In fact, for the Dicke model Dicke_1 ; Dicke_2 and the Lipkin-Meshkov-Glick model LMG_1 ; LMG_2 ; LMG_3 , the critical exponent satisfies $3/2$ as well, therefore the conventional SPT of the non-Hermitian cascade Rabi cavities we have studied belongs to the same universality class as those in the above mentioned systems.
@@ VII Summary
Herein, a non-Hermitian QRM consisting of two cascaded non-Hermitian Rabi cavities has been studied. Similar to Hermitian QRMs, the superradiance phase boundary of the non-Hermitian one can be exactly solved as well. Accompanied by the emergence of SPT, the ground state energy gets lower. Meanwhile, the atoms become active and radiating photons outward. However, the non-Hermiticity leads to the exotic phenomena without Hermitian counterpart. When the atom-optical field coupling phase satisfies $\varphi=\pm\frac{\pi}{4}$ or $\varphi=\pm\frac{3\pi}{4}$, the superradiance phase boundary is surprisingly dominated by a constant independent of the strength of the atom-cavity field coupling, suggesting the unconventional SPT. For other coupling phases, the superradiance phase boundary depends on the cavity coupling strength and the atom-optica field coupling strength together, suggesting the conventional SPT. Meanwhile, we have argued that the conventional SPT belongs to the second-order quantum phase transition, while the unconventional SPT belongs to the first-order quantum phase transition. Even though the model we have studied is not $\mathcal{PT}$-symmetric, there still exists fully real energy spectrum. By the photon population, atom population, and the fidelity we have known that there is no quantum criticality for the unconventional SPT due to characteristics of the absence of asymptotic process and the single-parameter sharp drop of fidelity. Differently, for conventional SPT, there exists quantum criticality, and the critical exponent extracted from the scaling analyze of fidelity susceptibility show that the SPT belong to the same universality class as the single Rabi model. With the development of circuit-QED experimental technique, the found SPTs may be realized in high-performance circuit- QED architecture.

### v2

@@ IV Signatures of the superradiance phase transition
To identify the signatures and establish the phase boundary of the Superradiance Phase Transition (SPT), we commence our investigation by analyzing the ground-state energy. Energies are ranked based on their real parts, with the ground-state energy denoting the lowest real part.
Under the condition of $\varphi=\frac{\pi}{2}$ and varying $\eta$, the real parts of the ground-state energies $\mathrm{Re}(E_{0})/\eta$ as a function of $\mathcal{J}$ for $g=0.7$ are depicted in Fig. 2(a). Here, the photonic truncation number $N=80$ ensures the convergence of results. It is observed that with the onset of the SPT, $\mathrm{Re}(E_{0})/\eta$ for different $\eta$ decreases from $-1$ to lower values. According to the analytical predictions, the transition point under $\varphi=\frac{\pi}{2}$ and $g=0.7$ is expected to occur at $\mathcal{J}=0.745$, marked by the vertical black line. This confirms the occurrence of the SPT at the theoretically derived transition points.
Similarly, the plot in Fig. 2(b) illustrates the relationship between $\mathrm{Re}(E_{0})/\eta$ and $\mathcal{J}$ for $\varphi=\frac{\pi}{4}$ with $g=0.4$. As in the $\varphi=\frac{\pi}{2}$ case, when $\mathcal{J}$ surpasses the transition point, $\mathrm{Re}(E_{0})/\eta$ for different $\eta$ decreases from $-1$ to lower values, indicating the onset of the SPT. The analytical phase boundary indicates that the transition point under $\varphi=\frac{\pi}{4}$ should be at $\mathcal{J}=0.5$, which aligns with the observed transitions for $g=0.4$.
The analysis of Fig. 1(c) shows that the second-order derivative of $\mathrm{Re}(E_{0})/\eta$ under $\varphi=\frac{\pi}{2}$ and $g=0.7$ is discontinuous at the transition point, characteristic of a second-order quantum phase transition. Conversely, Fig. 1(d) demonstrates that the first-order derivative of $\mathrm{Re}(E_{0})/\eta$ under $\varphi=\frac{\pi}{4}$ and $g=0.4$. These findings indicate that the conventional SPT is a second-order quantum phase transition, while the unconventional SPT is a first-order quantum phase transition.
Specifically, cases where $|\varphi|=\frac{\pi}{4}$ and $|\varphi|=\frac{3\pi}{4}$ correspond to second-order quantum phase transitions, whereas other values of $\varphi$ correspond to first-order quantum phase transitions. However, this study focuses primarily on conventional and unconventional superradiance phase transitions, thus excluding detailed consideration of results for other $\varphi$ values.
@@
In the superradiance phase, in addition to the macroscopic excitations of the photons, the atoms also become excited and active. To verify this, we plot the populations of atoms $\langle\sigma^{1(2)}_{z}\rangle$ as a function of $\mathcal{J}$ for different $\varphi$ and $g$ in Fig. 4(a)-4(d). When $\mathcal{J}$ is below the transition point, the atoms remain in their ground state with populations $\langle\sigma^{1(2)}_{z}\rangle=-1$. However, as $\mathcal{J}$ crosses the transition point, the populations deviate from -1, indicating that the atoms are excited and active. These excited atoms drive the macroscopic photon excitations observed in the SPT.
While the photon excitations and atom populations of the system change noticeably with the occurrence of SPT, the two types of SPTs exhibit distinct observable behaviors. As illustrated in Fig. 3, during a conventional SPT, $\langle a^{{\dagger}}_{1(2)}a_{1(2)}\rangle$ continuously varies to macroscopic values. In contrast, during an unconventional SPT, $\langle a^{{\dagger}}_{1(2)}a_{1(2)}\rangle$ abruptly jumps to macroscopic values without an asymptotic process. This distinction is similarly evident in the atomic populations shown in Fig. 4.
@@
In addition to the photon and atom populations, the excitation gaps associated with the two types of SPTs also exhibit distinct behaviors. These excitation gaps have two components: the real excitation gap ${\rm Re}(\Delta_{E})/\eta$, which is the gap between the real parts of the lowest two energies, and the imaginary excitation gap ${\rm Im}(\Delta_{E})/\eta$, which is the gap between the imaginary parts of the lowest two energies. Figures 5(a1)-5(d1) show ${\rm Re}(\Delta_{E})/\eta$ as a function of $\mathcal{J}$ for different $\eta$, while Figures 5(a2)-5(d2) present ${\rm Im}(\Delta_{E})/\eta$ under the same conditions.
Regardless of whether $\varphi=\frac{\pi}{2}$ or $\varphi=\frac{\pi}{4}$, the real excitation gap undergoes a closing process when $\mathcal{J}$ surpasses the transition point. However, a noticeable difference arises in the imaginary excitation gaps between the two cases. As seen in Figures 5(a2) and 5(b2), ${\rm Im}(\Delta_{E})/\eta$ remains zero (values much smaller than $10^{-14}$ can be numerically considered as zero) both before and after the transition point. In contrast, Figures 5(c2) and 5(d2) demonstrate that the imaginary excitation gaps experience a closing process similar to that of the real excitation gaps.
@@
For a $\mathcal{PT}$-symmetric Rabi system, a real-to-complex transition in the energy spectrum is well-established PT ; PT1 ; es_NH_2 ; es_NH_3 ; es_NH_4 . Remarkably, the non-Hermitian two-Rabi cavity system without $\mathcal{PT}$ symmetry also exhibits a real energy spectrum. Under the condition $\varphi=\frac{\pi}{2}$, the maximum imaginary part of the energy spectra for $g=0.7$ and $g=0.8$, denoted as ${\rm Im}(E)_{\text{max}}$, is plotted in Figs. 6(a) and 6(b), respectively. The analytically determined superradiance transition point is $\mathcal{J}=0.745$ for $g=0.7$ and $\mathcal{J}=0.82$ for $g=0.8$. ${\rm Im}(E)_{\text{max}}/\eta$ approaches zero both before and after the superradiance transition point, indicating that the energy spectra are real and no real-to-complex transition occurs.
In contrast, the spectral properties under $\varphi=\frac{\pi}{4}$ are distinct. Figures 6(c) and 6(d) show ${\rm Im}(E)_{\text{max}}/\eta$ for $g=0.4$ and $g=0.6$, respectively. Unlike the $\varphi=\frac{\pi}{2}$ case, the energy spectra under $\varphi=\frac{\pi}{4}$ are complex due to the non-zero ${\rm Im}(E)_{\text{max}}/\eta$. This distinction can be analytically understood from Eq. (10) and Eq. (13). When $\varphi=\frac{\pi}{2}$, the imaginary part vanishes, whereas for $\varphi=\frac{\pi}{4}$, it persists. The numerical results in Fig. 6 confirm these analytical insights.
The full real nature of the energy spectra for $\varphi=\frac{\pi}{2}$ suggests that the conventional SPT in this case shares the universality class with Hermitian QRM systems HP ; ps2 ; qc_1 ; qc_2 ; qc_3 ; zhangyy_1 , allowing for derivation of the critical exponent from its critical behavior. Conversely, for the complex energy spectrum case ($\varphi=\frac{\pi}{4}$), we hypothesize the absence of critical behavior. Although quantum criticality has been reflected in photon and atom populations for conventional SPT and its absence for unconventional SPT, we aim to validate these conjectures further through examination of ground-state fidelity and fidelity susceptibility.
@@
V Comparative Study of Conventional vs. Unconventional Superradiance Phase Transitions
We focus on the quantum phase transition induced by the cavity coupling parameter $\mathcal{J}$. The fidelity for the ground state $\mathcal{F}$ is defined as fidelity_0 ; fidelity_1 ; fidelity_2 ; fidelity_3 :
$\mathcal{F}=|\langle\psi_{0}(\mathcal{J})|\psi_{0}(\mathcal{J}+\delta\mathcal{J})|\rangle|,$ (14)
which measures the overlap of the ground state before and after a small change $\delta\mathcal{J}$ in $\mathcal{J}$. According to Refs. FS_1 ; FS_2 ; FS_3 ; FS_4 ; FS_5 , the fidelity susceptibility $\chi_{\mathcal{F}}$ is given by:
@@
If a quantum phase transition exists, $\mathcal{F}$ exhibits a drop at the transition point, while $\chi_{\mathcal{F}}$ peaks.
@@
Under $\varphi=\frac{\pi}{2}$, $\chi_{\mathcal{F}}/\eta$ for $g=0.7$ and $g=0.8$ are shown in Figs. 7(a) and 7(b), respectively. As $\mathcal{J}$ increases, $\chi_{\mathcal{F}}/\eta$ smoothly reaches a peak and subsequently decreases. Notably, $\chi_{\mathcal{F}}/\eta$ displays clear critical behavior; as $\eta$ increases, the peak at the coupling parameter $\mathcal{J}_{max}$ approaches the exact solution. Conversely, for $\varphi=\frac{\pi}{4}$, no critical phenomena are observed. As seen in Figs. 7(c) and 7(d), the fidelity curves sharply drop to zero at a single $\mathcal{J}$ value without any asymptotic behavior between the fidelity valley (or $\chi_{\mathcal{F}}$ peak) and $\eta$. This indicates the presence of quantum criticality in conventional SPT and its absence in unconventional SPT.
Next, we analyze the quantum criticality of conventional SPT ($\varphi=\frac{\pi}{2}$). From the $\chi_{\mathcal{F}}$ vs. $\eta$ curves, we extract the maximal fidelity susceptibility $\chi^{max}_{F}$ for a specific $\eta$. The universal finite $\eta$ scaling relation of $\chi_{\mathcal{F}}$ for $g=0.7$ is depicted in Fig. 8(a). The fitted black curve satisfies $\ln\chi_{\mathcal{F}}=(-1.363\pm 0.004)\ln\eta+2.388$, indicating a scaling exponent $\mu=-1.363\pm 0.004$. Following the relationship $\mu=\frac{2}{\nu}$ qc_1 ; RMP2011 , we derive $\nu=1.4674\pm 0.004$. Although numerically slightly deviating from the exact solution ($\nu_{exact}=3/2$) of the single Rabi model HP ; ps2 , we argue that the quantum phase transition under $\varphi=\frac{\pi}{2}$ shares the same universality class as the single Rabi model. Around the transition point, $\chi_{\mathcal{F}}$ scales as $\left[\chi^{max}_{\mathcal{F}}-\chi_{\mathcal{F}}\right]/\chi_{\mathcal{F}}=f\left[\eta^{1/\nu}\left(\mathcal{J}-\mathcal{J}_{max}\right)\right]$, where $f(x)$ is a universal scaling function. From Fig. 8(b), across various $\eta$, we observe that the data collapse into a single curve with the critical exponent $\nu=\frac{3}{2}$, indicating the same universality class as the single Rabi model HP ; ps2 . Indeed, similar critical exponents ($\nu=3/2$) are observed in the Dicke model Dicke_1 ; Dicke_2 and the Lipkin-Meshkov-Glick model LMG_1 ; LMG_2 ; LMG_3 , further affirming that the conventional SPT in the non-Hermitian cascade Rabi cavities studied here belongs to this universal class.
@@ VI Experimental feasibility
Similarly, experiments exploring our findings can also be conducted in ion trap setups. Ion traps utilize individual ions confined in harmonic traps as qubits, offering precise control over quantum states and interactions. This setup enables the emulation of phase-dependent interactions required to investigate PT-related physics, similar to circuit-QED architectures. The trapped-ion system offers a versatile platform combining laser-coupled spin and bosonic degrees of freedom, ideal for simulating Hamiltonians of light-matter interactions. Demonstrated capabilities include quantum simulations of many-body spin models experiment2 , single-ion quantum Rabi models experiment , and recently, the Rabi-Hubbard (RH) model with up to 16 ions experiment3 , exploring equilibrium phase transitions and quantum dynamical properties using spin observables. The scalability and controllability of ion trap systems make them suitable for exploring both conventional and unconventional superradiance phase transitions in non-Hermitian quantum Rabi-dimer models and other related phenomena.
@@ VII Summary
Exploring a non-Hermitian Quantum Rabi Model composed of two cascaded non-Hermitian Rabi cavities, we investigate its superradiance phase boundary akin to its Hermitian counterpart. As the system undergoes SPT, transitioning from a quantum vacuum state to one that becomes macroscopically populated, we observe unique exotic phenomena arising from non-Hermitian effects.
Notably, for specific atom-photon coupling phases such as $\varphi=\pm\frac{\pi}{4}$ or $\varphi=\pm\frac{3\pi}{4}$, the superradiance phase boundary remains constant, independent of the coupling strengths. This suggests an unconventional SPT. In contrast, for other coupling phases, the phase boundary depends on both the cavity and atom-photon coupling strengths, indicative of a conventional SPT.
We argue that the conventional SPT corresponds to a second-order quantum phase transition, whereas the unconventional SPT exhibits characteristics of a first-order quantum phase transition. Despite the model’s non-$\mathcal{PT}$-symmetry, we observe a fully real energy spectrum under certain conditions.
Analysis of photon and atom populations, as well as fidelity measurements, indicates that the unconventional SPT lacks quantum criticality, characterized by the absence of asymptotic behavior and a sharp drop in fidelity at a single parameter value. In contrast, the conventional SPT exhibits quantum criticality, confirmed by critical exponents extracted from fidelity susceptibility scaling analyses, showing agreement with the single Rabi model’s universality class.
Recent advancements in circuit-QED and ion trap experimental techniques offer substantial potential to implement and verify these identified SPTs in high-performance, state-of-the-art architectures. Moreover, these capabilities can be extended to investigate more complex coupled light-matter systems that currently surpass the capabilities of classical computational simulations.